Answers 2b is not 9000 and part 3 is not 335000. Please provide work.Determining the Optimal Product Mix with One Constrained Resource and a Sales Constraint Comfy Fit Company manufactures two t
Answer #1

2

Optimal mix:

First preference is given to Swoop.

Maximum production of swoop is 40,000 units and thus, 40,000 units of swoop are produced.

Balance hours left of machine = 1,000 hours X 7 - 40,000 X 6 minutes/ 60 minutes = 3,000 hours

Machine time per rufus = 20/60 = 0.33 hours

Rufus production = 3,000 hours / 0.33 hours = 9,091 units

3

Contribution = 40,000 Swoop X $5 + 9,091 Rufus X $15 = 336,365